Product Cost Method of Product Costing

Voice Com, Inc., uses the product cost method of applying the cost-plus approach to product pricing. The costs of producing and selling 5,030 units of cell phones are as follows:

Variable costs: Fixed costs:
Direct materials $88 per unit Factory overhead $201,800
Direct labor 35 Selling and admin. exp. 71,500
Factory overhead 26
Selling and admin. exp. 19
Total variable cost per unit $168 per unit

Voice Com desires a profit equal to a 16% rate of return on invested assets of $601,600.

a. Determine the amount of desired profit from the production and sale of 5,030 units of cell phones.
$

b. Determine the product cost per unit for the production of 5,030 of cell phones. If required, round your answer to nearest dollar.
$ per unit

c. Determine the product cost markup percentage (rounded to two decimal places) for cell phones.
%

d. Determine the selling price of cell phones. Round to the nearest dollar.

Total Cost $per unit
Markup per unit
Selling price $per unit
Answer #1

a.

Desired profit = Invested assets * Rate of return on invested assets

= 601,600 * 16%

= 96,256

b.

Total fixed costs per unit = (Factory overhead + Selling and administrative exp.) / 5,030 units

= (201,800+71,500) / 5,030

= 54.33

Product cost per unit = Total varaible cost per unit + Total fixed cost per unit

= 168 + 54.33

= 222.33

c.

Markup percentage = Desired profit / Total costs

= 96,256 / [(168*5,030) + 201,800 + 71,500]

= 96,256 / 1,118,340

= 8.61%

d.

Total cost 222.33
Markup 19.14 (222.33*8.61%)
Selling price 241.47